OpenStudy (anonymous):

Find the domain, intercepts, the intervals where function is increasing and decreasing, extreme values, intervals of concavity, points of inflection, asymptotes (vertical, horizontal, or slant). Show the1st and the 2nd derivative tables. Sketch a graph. (x^2 +12)/(x-1)

OpenStudy (anonymous):

hmmm... looks familiar...

OpenStudy (anonymous):

haha yeah after we stopped i was totally lost

OpenStudy (anonymous):

and now i am like totally stressing because all this is due in 2 hours and i dont know if i can finish

OpenStudy (anonymous):

\[\huge f(x)=\frac{x^2+12}{x-1} \] \[\huge f'(x)=\frac{x^2-2x-12}{(x-1)^2} \] i guess that's where we left off...

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

let's do all the things that don't require calculus first...domain, intercepts, asymptotes...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

look at f(x)... is there any x where f(x) is NOT defined?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

ok... good... do domain = {all real except x=1} this by the way is also where your vertical asymptote occurs... at x=1 is your vertical asymptote.

OpenStudy (anonymous):

in a rational function, whatever makes the denominator = 0 is a vertical asymptote.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now let's look for x and y intercepts...

OpenStudy (zzr0ck3r):

lol why new question?

OpenStudy (anonymous):

because I need as much help as a can get! ha

OpenStudy (zzr0ck3r):

but it was sill open? I wrote a bunch of stuff for you.

OpenStudy (anonymous):

to find the x-intercept, set y=0 so you need to solve: \[\huge 0=\frac{x^2+12}{x-1} \]

OpenStudy (anonymous):

what it was? im sorry i will try to go find it and look at that along with this, i dont know how to use this website very well

OpenStudy (anonymous):

do you have a solution to this equation?

OpenStudy (anonymous):

square root of -1/12?

OpenStudy (anonymous):

i think you mean sqrt(-12) ?

OpenStudy (anonymous):

if you say so :) i was kind of confused on how to solve that one

OpenStudy (anonymous):

the only thing that matter is that numerator... if that numerator is 0, then it doesn't matter what the denominator is because you'll have 0 over some number... which is zero...

OpenStudy (anonymous):

so you're actually solving this: \(\large 0=x^2+12 \)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now, is the solution to that equation a real number?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so what does that tell you about the x-intercept(s)?

OpenStudy (anonymous):

there arent any?

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

now let's look for the y-intercept..

OpenStudy (anonymous):

what do i do to find that?

OpenStudy (anonymous):

just set x=0 and figure it out f(0)

OpenStudy (anonymous):

using the equation from earlier?

OpenStudy (anonymous):

the function...

OpenStudy (anonymous):

oh oh ok

OpenStudy (anonymous):

-12?

OpenStudy (anonymous):

so your y-intercept is at y=-12...

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

ok.. 1 more precalculus thing before we do the calculus...

OpenStudy (anonymous):

\[\huge f(x)=\frac{x^2+12}{x-1} \] in precalculus if the degree of the numerator = the degree of the denominator that you have a horizontal asymptote... do you have one here?

OpenStudy (anonymous):

the numerator is greater than the demonimator

OpenStudy (anonymous):

so do we have a horizontal asymptote?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if the degree of the numerator is 1 more than the degree of the denominator, then you have a slant asymptote.... as you can see, we have a slant asymptote... agreed?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

do you know how to get the equation of the slat asymptote?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

have you ever worked with slant asymptotes?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

that i can remeber anyways

OpenStudy (anonymous):

to get the equation of the slant asymptote, all you have to do is do LONG DIVISION..... YAY!!!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):